### v1

# Visatronic: A Multimodal Decoder-Only Model for Speech Synthesis

Abstract
In this paper, we propose a new task – generating speech from videos of people and their transcripts (VTTS) – to motivate new techniques for multimodal speech generation. This task generalizes the task of generating speech from cropped lip videos, and is also more complicated than the task of generating generic audio clips (e.g., dog barking) from videos and text. Multilingual versions of the task could lead to new techniques for cross-lingual dubbing. We also present a decoder-only multimodal model for this task, which we call Visatronic. This model embeds vision, text and speech directly into the common subspace of a transformer model and uses an autoregressive loss to learn a generative model of discretized mel-spectrograms conditioned on speaker videos and transcripts of their speech. By embedding all modalities into a common subspace, Visatronic can achieve improved results over models that use only text or video as input. Further, it presents a much simpler approach for multimodal speech generation compared to prevailing approaches which rely on lip-detectors and complicated architectures to fuse modalities while producing better results. Since the model is flexible enough to accommodate different ways of ordering inputs as a sequence, we carefully explore different strategies to better understand the best way to propagate information to the generative steps. To facilitate further research on VTTS, we will release (i) our code, (ii) clean transcriptions for the large-scale VoxCeleb2 dataset, and (iii) a standardized evaluation protocol for VTTS incorporating both objective and subjective metrics.
1 Introduction
Figure 1: Visatronic overview. In addition to existing text-to-speech (left top) and lips-to-speech tasks (right top), we propose a novel multimodal generative task (bottom), video-text-to-speech (VTTS), where the model is conditioned on the video of talking people and corresponding text transcriptions in order to generate speech. Also, we propose a unified multimodal decoder-only architecture, Visatronic, that processes all modalities (video $\mathbf{v}$ (grey), text $\mathbf{t}$ (grey), and speech $\mathbf{s}$ (blue)) in the LM-style transformer model after all modalities are discretized. The model is trained using cross entropy loss $\mathcal{L}_{CE}$ computed on speech discrete values $\mathbf{s}_{t}$ given the mixed multimodal input sequence ${\{\mathbf{v}\}_{t},\{\mathbf{t}\}_{i},\{\mathbf{s}\}_{t}}$. Each input modality is processed in the unified framework, enabling the model to learn interactions between different modalities while learning the temporal alignment.
The research community has made strides in building multimodal models for speech and audio generation. These techniques have been driven by two different types of problems: {NoHyper} **footnotetext: Work done during internship at Apple. generating speech from cropped videos of lips [46], and generating audio (e.g. barking of dogs) from textual descriptions and videos [19]. The former problem simplifies the task of video-conditioned speech generation by using a pretrained model to crop out lips, while the latter deals with generating outputs whose content is very loosely specified and does not need to show as strong correspondence as speech does to a text sequence. In this paper, we propose a new task – generating speech from videos of people and their transcripts (VTTS) – to motivate new techniques for multimodal speech generation. VTTS is more complicated than the above tasks in several different ways. Firstly, the task is defined as an end-to-end task, in that it does not require additional models to detect and crop the lips in the videos. Secondly, the synthesis must satisfy multiple critical criteria: the speech must be clearly intelligible by following the input text, precisely synchronized with the speaker’s movements, and sound natural in terms of prosody and speaking style. In addition, it should leverage facial features that are informative to the task of speech generation, such as emotion and intensity, and also be consistent with other events in the video. We believe VTTS can enable novel applications beyond existing speech generation tasks. For example, multilingual models trained with this approach could be used to perform video dubbing across different languages.
Multimodal generative modeling has made rapid strides recently using auto-regressive transformer models [19, 48, 26] and can be applied to VTTS. These methods piggyback on the observation that transformer-based large language models (LLMs) can learn extremely complicated distributions using next-step prediction. In order to do so, these approaches typically use a vector-quantized variational autoencoder (VQ-VAE) [40] to convert the inputs from the different modalities to sequence of discrete tokens that the language model can consume. Using this recipe, prior work has been able to generate data such as videos, images and speech conditioned on text input [48, 19, 45, 4].
Recently, it has also been shown that a similar autoregressive approach can be used with joint models of text and speech, without tokenization by simply quantizing the mel-spectrogram of speech into discrete, uniformly spaced bins [3]. In this paper, we show that this approach can be generalized and applied to VTTS. We call our model Visatronic. Visatronic embeds each of the modalities – text, vision and speech – into the embedding space of the transformer. Text is input to the model by tokenization followed by embedding lookup. Videos are converted to discrete representations using a VQ-VAE and Visatronic learns to embed them through a special embedding scheme. Speech is quantized and embedded through a similar scheme to the vision inputs.
To evaluate Visatronic’s effectiveness in real-world scenarios, we conduct extensive experiments on the LRS3 [1] dataset following [46], and a more challenging dataset VoxCeleb2 [9], which contains “in-the-wild” videos featuring hundreds of unique speakers with unconstrained vocabulary and diverse acoustic conditions. Compared to LRS3, VoxCeleb2 is 3x larger and contains paired video-speech data without text, has a more diverse and larger pool of speakers and acoustic conditions, and larger background noise. Given these factors, we mainly focus on the VoxCeleb2 dataset in the paper.
To the best of our knowledge, there are no standarized evaluation protocols for VTTS, so we establish a comprehensive evaluation framework that combines both subjective human assessments and objective metrics. Furthermore, we implement and evaluate multiple baseline approaches to provide meaningful comparisons and facilitate future research in this emerging field. Our results demonstrate that Visatronic performs better than prior techniques that use either cropped lips or text as inputs, achieving 12.2% word error rate (WER) on VoxCeleb2 [9] and 4.5% WER on LRS3 [1] datasets. These results also demonstrate that Visatronic generalizes robustly to diverse visual and acoustic conditions not seen during training.
Our contributions are summarized below as follows:
  • We propose a new multimodal generative task, video-text-to-speech (VTTS), to facilitate research in multimodal generation and understand importance of the video conditioning for speech generation.
  • We show the importance of the data processing pipeline to prepare triplets of (video, text, speech) for model training and provide the clean transcriptions for the VoxCeleb2 [9].
  • We successfully trained a unified multimodal decoder-only model for speech generation. We show that conditioning on both video and text improves speech generation over the TTS models across both objective and subjective metrics, e.g. word error rate of a speech recognition model on the generated speech is reduced by more than relative 15%.
  • We formulate an evaluation protocol for VTTS, that incorporates existing objective and subjective metrics, and defines a new objective metric, TimeSync, to measure time alignment between generated and ground truth speech.
2 Visatronic
Figure 2: Video representation. The input video frame at time $t$ is mapped via encoder of VQ-VAE model to the downsampled spatial representation being in $\mathbb{R}^{H^{\prime}\times W^{\prime}\times D}$. Then every spatial location $(h,w)$, with a vector representation in $\mathbb{R}^{D}$, e.g. $(0.8,\dots)$, is mapped to a discrete value, e.g. $3$, using the learned VQ-VAE codebook $\mathbf{C}^{v}$ by finding the closest codebook element via $l_{2}$ similarity. The discrete value at spatial location $(h,w)$ and time $t$ is mapped to a representation $\mathbb{R}^{D^{\prime}}$ via a learnable embedding layer, where $D^{\prime}$ is the transformer decoder dimension. Finally, we perform different aggregations across spatial grid $H^{\prime}\times W^{\prime}$ for embeddings to obtain the final embedding for the frame before inputting it to the transformer decoder: (a) a self-attention mechanism where query $Q$, key $K$, and value $V$ transformations learn spatial relationships within the region, (b) summation or (c) mean pooling across all embeddings to capture aggregate spatial information, (d) max pooling across all embeddings to capture the most salient features, (e) stacking all embeddings followed by a learnable linear projection $\mathbb{R}^{H^{\prime}W^{\prime}D^{\prime}}\to\mathbb{R}^{D^{\prime}}$.
In the rest of the paper, we denote tensors as $\mathbf{x}$ while $\mathbf{x}_{i,...}$ denotes the $(i,...)$-th component of tensor $\mathbf{x}$.
2.1 Video-Text-To-Speech (VTTS)
Video-text-to-speech synthesis (VTTS) can be formulated as follows: given (a) the input video frames of the speaker $\mathbf{x}^{v}\in\mathbb{R}^{T^{v}\times H\times W\times 3}$, where $H$ and $W$ are the spatial video resolution (frame height and width, respectively) and $T^{v}$ is the total number of frames in the video; and (b) text tokens $\{\mathbf{x}^{t}_{i}\}_{1}^{N}$ representing the transcript of speech in the video, where $\mathbf{x}^{t}_{i}\in\textit{Vocabulary}$, $N$ is length of the tokenized transcript, the goal is to generate speech signal $\mathbf{x}^{s}\in\mathbb{R}^{T^{s}}$, where $T^{s}$ is length of speech signal, such that spoken words in speech correspond to the written text $\{\mathbf{x}^{t}_{i}\}_{1}^{N}$, and video and speech are aligned in time.
2.2 Input Representation
Video Representation
To obtain a latent representation of the video input $\mathbf{x}^{v}$, we leverage a pretrained VQ-VAE model [45] which is pre-trained on the general video dataset, Kinetics-600  [5], with codebook $\mathbf{C}^{v}=\{\mathbf{c}_{1}^{v},\mathbf{c}_{2}^{v},\dots,\mathbf{c}_{K^{v}}^{v}\}$ of size $|\mathbf{C}^{v}|=K^{v}$, where $\mathbf{c}_{i}\in\mathbb{R}^{D}$. Using the encoder of the VQ-VAE, we map each input video frame $\mathbf{x}^{v}_{t}\in\mathbb{R}^{H\times W\times 3}$ to a latent vector $\mathbf{y}^{v}_{t}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times D}$ with downsampled resolution $H^{\prime}\times W^{\prime}$. Each spatial element in $\mathbf{y}^{v}_{t}$ is then mapped to the index of its nearest codebook entry $\mathbf{v}_{t,h,w}\in\mathbb{C}^{v}=\{1,\cdots,K^{v}\}$ based on $\ell_{2}$ distance. Thus, every input video frame $\mathbf{x}^{v}_{t}$ is represented as $\mathbf{v}_{t}\in[{\mathbb{C}^{v}}]^{H^{\prime}\times W^{\prime}}$ – a set of discrete values, see Figure 2.
We use the VQ-VAE model to discretize video due to its ability to compress the video representation while preserving both spatial and temporal dynamics crucial for video understanding. Concretely, the pre-trained VQ-VAE model [45] compresses videos with $H\times W=224\times 224$ spatial resolution to $H^{\prime}\times W^{\prime}=16\times 16$ spatial resolution with the codebook dimension $D=3264$ and codebook size $K^{v}=2048$. Although this VQ-VAE model is pre-trained on the general videos, we found it reconstructs speakers videos with sufficient quality to preserve necessary spatial information, see Section D in Appendix.
Figure 3: Speech representation. We follow the speech discretization process from dMel [3]: each continuous mel-filterbank at time $t$ extracted from the raw audio is mapped into a discrete values using a codebook of evenly spaced values in the range $[m,M]$ by taking the closest codebook value, where $m$ and $M$ are the minimum and maximum values of log mel-filterbanks computed across the dataset. Afterwards, each discretized log mel-filterbank at time $t$ is mapped through a learnable embedding layer, all representations for log mel-filterbanks at time $t$ are stacked together and resulting vector is projected by a learnable linear layer to the model dimention $D^{\prime}$. The example illustrates this process by converting log mel-filterbank values $(5.1,2.8,-0.4,\dots)$ into bin indices $(10,8,2,\dots)$, which are then embedded for the Visatronic model processing. Note, that all discretized log mel-filterbanks at time $t$ are predicted in parallel and independently by a decoder-only model.
Following quantization, every discrete value is mapped via a learnable embedding layer $\mathbf{E}^{v}(\cdot):\mathbb{C}^{v}\to\mathbb{R}^{D^{\prime}}$ to $\mathbf{e}^{v}_{t,h,w}$. The representation for the whole frame after embedding is $\mathbf{e}_{t}^{v}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times D^{\prime}}$, where $D^{\prime}$ is the Visatronic transformer input dimension. Subsequently, we explore various methods for aggregating the spatial dimensions of this representation prior to inputting to the transformer decoder:
Attention: having learnable $Q,K,V\in\mathbb{R}^{D^{\prime}\times D^{\prime}}$ and $\text{attn}_{h,w}=\text{softmax}_{h,w}(Q\mathbf{e}^{v}_{t,1,1},K\mathbf{e}^{v}_{t,h,w})$ we compute $\mathbf{z}_{t}^{v}=\frac{1}{\sqrt{D^{\prime}}}\sum_{h=1}^{H^{\prime}}\sum_{w=1}^{W^{\prime}}\text{attn}_{h,w}\,\,\,V\mathbf{e}^{v}_{t,h,w}$;
Summation: $\mathbf{z}_{t}^{v}=\sum_{h=1}^{H^{\prime}}\sum_{w=1}^{W^{\prime}}\mathbf{e}^{v}_{t,h,w}$;
Mean pooling: $\mathbf{z}_{t}^{v}=\frac{1}{H^{\prime}W^{\prime}}\sum_{h=1}^{H^{\prime}}\sum_{w=1}^{W^{\prime}}\mathbf{e}^{v}_{t,h,w}$;
Max pooling: $\mathbf{z}_{t}^{v}=\max_{(h,w)}\mathbf{e}^{v}_{t,h,w}$;
Stacking: stack embeddings and then project it via a learnable linear layer $\mathbf{L}^{v}(\cdot):\mathbb{R}^{H^{\prime}W^{\prime}D^{\prime}}\to\mathbb{R}^{D^{\prime}}$,
$\mathbf{z}_{t}^{v}=\mathbf{L}^{v}([\mathbf{e}^{v}_{t,1,1},\mathbf{e}^{v}_{t,1,2},\dots,\mathbf{e}^{v}_{t,1,W^{\prime}},\mathbf{e}^{v}_{t,2,1},\dots,\mathbf{e}^{v}_{t,H^{\prime},W^{\prime}}])$.
As we show later, this multi-faceted approach enables effective capture of both local and global video characteristics in Visatronic.
Text Representation    For text processing, we employ a character-level tokenizer that maps the input text $\{\mathbf{x}^{t}_{i}\}_{1}^{N}$ to a sequence of discrete tokens $\mathbf{t}_{j}\in\mathbb{C}^{t}=\{1,2,\dots,K^{t}\}$ with $|\mathbb{C}^{t}|=K^{t}$, followed by a learnable embedding layer $\mathbf{E}^{t}(\cdot):\mathbb{C}^{t}\to\mathbb{R}^{D^{\prime}}$. Character-level tokenization reduces vocabulary size $K^{t}$ and improves generalization by capturing fine-grained linguistic features.
Speaker Representation    For multi-speaker modeling, we extract speaker representations using a pre-trained dvector model [41] that produces 512-dimensional embeddings. These speaker embeddings are then projected through a learnable linear layer to match the model dimension $D^{\prime}$. The speaker embeddings are required for this task to maintain speaker characteristics.
Speech Representation    We utilize dMel [3], a simple yet effective discretization approach for speech processing; see Figure 3 for an overview. Given an input speech signal $\mathbf{x}^{s}$, we first compute continuous log mel-filterbanks $\mathbf{y}_{t}^{s}\in\mathbb{R}^{F}$ for a frame at time $t$, where $F$ is the number of log mel-filterbanks. Then, we map every log mel-filterbank $\mathbf{y}_{t,f}^{s}\in\mathbb{R}$ to a discrete value $\mathbf{s}_{t,f}\in\mathbb{C}^{s}=\{1,2,\dots 2^{K^{s}}\}$ using a codebook $\mathbf{C}^{s}=\{\mathbf{c}^{s}_{1},\mathbf{c}^{s}_{2},\dots,\mathbf{c}^{s}_{2^{K^{s}}}\}$: $\mathbf{c}^{s}_{i}\in\mathbb{R}$ are evenly spaced values in the range $[m,M]$, where $m$ and $M$ are the minimum and maximum values of mel-filterbanks computed across the dataset. To discretize, we take the closest codebook value, i.e., $\mathbf{s}_{t,f}=\text{argmin}_{i\in\mathbb{C}^{s}}|\mathbf{y}_{t,f}^{s}-\mathbf{c}^{s}_{i}|$.
After each speech frame is discretized, every discrete value is mapped via a learnable embedding layer $\mathbf{E}^{s}(\cdot):\mathbb{C}^{s}\to\mathbb{R}^{d^{\prime}}$ to a representation $\mathbf{e}^{s}_{t,f}$. The representation for the whole frame is given by $\mathbf{e}_{t}^{s}\in\mathbb{R}^{F\times d^{\prime}}$, where $d^{\prime}$ is the intermediate dimension. Subsequently, we stack these embeddings and project the resulting vector to a final embedding $\mathbf{z}^{s}_{t}\in\mathbb{R}^{D^{\prime}}$ via a learnable linear layer $\mathbf{L}^{s}(\cdot):\mathbb{R}^{Fd^{\prime}}\to\mathbb{R}^{D^{\prime}}$: $\mathbf{z}_{t}^{s}=\mathbf{L}^{s}([\mathbf{e}^{s}_{t,1},\mathbf{e}^{s}_{t,2},\dots,\mathbf{e}^{s}_{t,F}])$.
This training-free discretization enables effective processing of speech signals in our framework. Following [3] we use $K^{s}=4$ bits with $|\mathbb{C}^{s}|=16$, $F=80$ log mel-filterbank channels and $d^{\prime}=24$.
Speech Inversion    To reconstruct the speech signal $\mathbf{x}^{s}$ from the speech discrete values $\mathbf{s}_{t,f}$ predicted by the multimodal transformer decoder (Section 2.3), we follow [3]: first, we transform the indices back to the log mel-filterbanks via the codebook $\mathbf{C}^{s}$: $\mathbf{\hat{y}}^{s}_{t,f}=\mathbf{c}^{s}_{\mathbf{s}_{t,f}}$. Subsequently, we apply a vocoder [44] to transform reconstructed log mel-filterbanks $\mathbf{\hat{y}}^{s}_{t,f}$ back into the time domain signal $\mathbf{x}^{s}$. The vocoder is trained independently and is not part of the Visatronic transformer decoder-based model.
2.3 Unified Multimodal Video-Text-Speech Transformer Decoder
We propose a unified multimodal decoder-only transformer architecture for processing multiple modalities – video, text and speech – in order to generate speech given video and text inputs, see Figure 1. The architecture consists of a single transformer decoder that processes the multimodal input representations from Section 2.2. Unlike traditional approaches that use one modality as input, or separate encoder(s) for multimodal input, our unified architecture enables cross-modal interactions through self-attention layers while maintaining temporal coherence. The model is trained end-to-end using cross entropy loss to predict the next discrete values in sequence, allowing it to learn intrinsic relationships across modalities that are crucial for tasks requiring multimodal understanding. During inference, the model can generate tokens autoregressively while maintaining coherence across all modalities.
Integration of Multimodal Sequences    For effective processing of multiple modalities with different temporal resolutions, we implement various input mixing strategies, see Figure 4. The fundamental challenge lies in handling different sampling rates and temporal ordering: speech inputs from dMel are sampled at 25ms intervals $(0.00s,0.025s,0.05s,\dots)$, whereas 25fps video inputs are sampled at 40ms intervals $(0.00s,0.04s,0.08s,\dots)$, and text tokens appear sparsely in the sequence. We explore the following ways to combine different modalities’ inputs into one sequence:
  • Ordering Strategy: Representations from all modalities are temporally ordered: either text, video and speech inputs; or video, text and speech inputs. For both cases, when speech is generated, the transformer decoder attends to all representations of text and video modalities. The ordering between text and video defines the interplay between them.
  • Streaming Strategy: Text tokens go first, but video and speech inputs are ordered following their original time alignment, thus preserving the natural flow of information in each of these modalities. In this approach, the speech inputs never attend to the future in time video inputs, thus reducing the sequence length processed at every speech generation step.
Positional Encoding    Due to combination of both video and text modalities for speech generation, our sequences are longer than in TTS task. Thus, capturing positional information properly is crucial. Prior work consistently showed that relative positional embeddings perform better (see e.g. [38, 3]). We apply RoPE [37] multiplicative relative positional embedding across the entire sequence. As a simple way, we maintain a global position space across all modalities, treating speaker, video, text and speech inputs uniformly in terms of positional embedding. In addition, thanks to video and speech time alignment, we investigate different position sequences to align representations appearing at the similar timestamps from different modalities, see positions notation in Figure 4.
Initialization    Placing all modalities’ inputs into one sequence for the decoder, we found that having different submodules to map each modality to the shared space leads to inconsistency of the embeddings across modalities (e.g. they have very different norm magnitudes). Thus, proper initialization of these submodules is essential. We identified that a proper scale for the initial weights distribution by bringing all inputs’ final embeddings to the same sphere is sufficient for stable and fast convergence during training.
Robust Training    Our unified decoder model is trained to predict speech discrete representations while being conditioned on all modalities during inference. During training we compute the cross-entropy loss $\mathcal{L}_{CE}$ only on the speech discrete representations, omitting the loss on others. All $F$ discrete log mel-filterbanks at each timestamp $t$ are predicted independently and in parallel. To ensure robust training, we follow dMel training observations and apply random span masking with probability $p$ to video, text and speech representations, forcing the model to leverage cross-modal information rather than relying solely on one modality. Speech masked regions are excluded from the loss computation. During inference, the model autoregressively generates speech discrete representations while being conditioned on speaker information, video and text.
Figure 4: Input sequence layout for Visatronic. We encode all the modalities to a discrete space (see Figures 2 and 3) directly used by the decoder transformer model: each modality discrete representations are marked with its colored square. Each row represents the different strategies we adopt for combining multimodal information for learning the temporal alignment between these modalities: (top) text goes before video, while video is followed by speech; (middle) text goes first, while speech and video are ordered in time so that speech generation at time $t$ attends to the whole text but only past video at $t^{\prime}<t$; (bottom) video goes before text while text if followed by speech. Positions sequence either has global indexing across all modalities, or video and speech are aligned in time.
3 Experiments
Method Input Modality GT WER ($\downarrow$) GT (discrete) WER ($\downarrow$) WER ($\downarrow$) Sync Score ($\uparrow$) TimeSync (s) ($\downarrow$)
TTS Text 4.0 ±0.1 10.5 ±0.1 19.0$\mathord{+}8.5$
VTTS (VT-ordered) Video-Text 17.2$\mathord{+}6.7$
TTS Text 2.6 ±0.1 10.1 ±0.2 14.7$\mathord{+}4.7$ 1.54 0.62 ±0.98
[…]
VTTS (VT-ordered) Video-Text $\mathord{+}2.1$ 1.64 0.47 ±0.63
Table 1: Word error rate (WER) and TimeSync on VoxCeleb2. We report ground truth (GT) WER (computed on original audio), GT (discrete) WER (computed on the reconstructed audio from ground truth discrete speech tokens), and WER computed on generated speech. WER is calculated between ground truth text and transcription obtained by forwarding whisper-large v2 on selected audio. The first set of results (first two rows) uses PL.v1 transcriptions, while the remaining results are on PL.v2. VTTS (VT-ordered) achieves the best performance with 12.2% WER, outperforming both single-modality (TTS) and other multi-modal approaches. GT (discrete) WER (10.1±0.2%) represents the theoretical lower bound by excluding the performance loss due to speech discretization. Our goal is to minimize the gap between model WER and GT (discrete) WER, with VTTS (VT-ordered) achieving closest alignment to this lower bound with $\sim$2.1difference from its GT (discrete) WER. TimeSync shows that the video modality provides better synchronization between video and generated speech.
Lip2Speech† [18] SVTS† [29] VCA-GAN† [17] DiffV2S† [7] LipVoicer† [46] VTTS (TV-ordered) VTTS (VT-ordered)
WER 57.4 82.4 90.6 39.2 21.4 4.5 8.2
Table 2: Generalization capability on LRS3. Word Error Rate (WER) results demonstrate strong generalization ability of our models trained on VoxCeleb2 when evaluated on LRS3 without any training on LRS3. Both VTTS (TV-ordered) and VTTS (VT-ordered) variants significantly outperform existing methods that were specifically trained on LRS3 (denoted by †).
Datasets    1) LRS3 [1] is audio-visual dataset in English, compiled from TED and TEDx video presentations. This dataset stands out for its focus on unconstrained long sentences, featuring a rich vocabulary of over 50k words and thousands of unique speakers. It contains approximately 151k videos with around 439h of speech with transcription. There are 1,452 videos in the test split. 2) VoxCeleb2 [9] is a large-scale audio-visual dataset primarily designed for speaker recognition task but applicable to various audio-visual processing domains. It consists of over 1M face-cropped YouTube videos from more than 6k distinct identities, resulting in 1.6k hours of speech w/o paired transcription. The dataset is characterized by high variability in lighting conditions, image quality, pose, and motion blur, with an average video duration of 8s. This diversity in real-world conditions makes VoxCeleb2 particularly useful for developing robust models capable of performing well in unconstrained environments. To train our models on VoxCeleb2, we first develop a pipeline for pseudo-labeling (PL) the speech using Demucs [10] for speech enhancement, Whisper-large v2 [33] for automatic transcription, and proper data filtering as the data are multilingual. The initial version of labeled data, PL.v1, was obtained by keeping English-only detected samples. Later, we improved upon it by additional filtering of inconsistent too long or too short transcriptions, leaving us with PL.v2 version of data. To evaluate our models, we randomly selected subset of 2k samples from the test set.
Objective Evaluation Metrics.    To evaluate how well generated speech preserves content information, we use the word error rate (WER) metric computed between the speech recognition model outputs from Whisper-large v2 on the audio samples and the ground truth transcripts. The synchronization score (SyncScore) is computed using the pre-trained model from [8]. This model is trained to predict the time-offset between lip crops and audio based on the distance between visual and audio embeddings over a sliding window of frames. The confidence score is computed as the difference between the median and minimum distances over this sliding window and was originally used to determine the active speaker in a multi-speaker video. From evaluation, we found that SyncScore fails in many cases and does not properly measure TTS model synchronization. For that reason, we propose a new metric, TimeSync: we take ground truth transcription and do force alignment of its phoneme sequence to each audio via HMM model from HTK [47]; the latter gives us location in time for each phoneme; finally, we compute the average absolute time difference between locations of centers of the phoneme segments for ground truth and generated audio and average across all phonemes in the test set.
Subjective Evaluation Metrics.    We randomly selected 50 samples each from VoxCeleb2 and LRS3 test data for human evaluation to assess the naturalness, intelligibility and synchronization of the generated speech following [46]. Using Mean Opinion Score (MOS) with 95% confidence intervals, human evaluators rated the speech naturalness, intelligibility and synchronization on a scale of 1 to 5, where 1 represents the worst and 5 the best quality. Details on the full protocol are provided in Appendix, Section E.
Implementation Details.   For implementation details and training configuration, refer to Appendix, Section F.
3.1 State-of-the-art Comparison
Table 1 shows a comparison of our proposed approaches and the TTS baseline trained and evaluated on VoxCeleb2 data. All results show that video brings improvement into both content generation and time synchronization.
We further evaluate on LRS3 data models which were trained on VoxCeleb2 data only. Results are shown in Table 2: VTTS (TV-ordered) achieves 4.5% WER, surpassing even LipVoicer’s 21.4% WER by a large margin, while maintaining a small gap of 2.1% from its GT (discrete) WER lower bound. These results demonstrate our models’ robust generalization to out-of-distribution data and different speaking conditions.
3.2 Human Evaluation Results
Human evaluation, presented in Tables 3 and 4, shows that VTTS (VT-ordered) achieves the best performance in Intelligibility (3.48) and Naturalness (3.20), while VTTS (TV-ordered) performs better in Synchronization (2.50). These scores approach the GT (discrete) upper bound, demonstrating the effectiveness of our proposed variants.
3.3 Ablations
Faster convergence   Table 5 shows results for different number of training steps. Our models achieve better performance at 2M iterations compared to TTS baseline, showing that video modality is speeding up training convergence.
Method Intelligibility ($\uparrow$) Naturalness ($\uparrow$) Synchronization ($\uparrow$)
GT 4.55 ±0.09 4.79 ±0.05 4.57 ±0.10
GT (discrete) 3.95 ±0.13 3.77 ±0.15 4.36 ±0.12
TTS 3.17 ±0.19 2.92 ±0.21 1.98 ±0.15
VTTS (TV-streaming) 3.19 ±0.17 2.99 ±0.16 2.28 ±0.17
VTTS (TV-ordered) 3.35 ±0.17 3.02 ±0.19 2.50 ±0.21
VTTS (VT-ordered) 3.48 ±0.15 3.20 ±0.19 2.48 ±0.19
Table 3: Human evaluation on VoxCeleb2. Mean Opinion Scores (MOS) (1-5) with 95% confidence intervals for Intelligibility, Naturalness, and Synchronization metrics. VTTS (VT-ordered) achieves the best performance in Intelligibility (3.48) and Naturalness (3.20), while VTTS (TV-ordered) performs best in Synchronization (2.50). Both significantly outperform TTS baseline across all metrics. Ground truth (GT) is an upper bound, while GT (discrete) is an upper bound due to speech discretization.
Method Intelligibility ($\uparrow$) Naturalness ($\uparrow$) Synchronization ($\uparrow$)
GT 4.79 ±0.05 4.79 ±0.05 4.73 ±0.06
GT (discrete) 4.32 ±0.11 3.80 ±0.11 4.59 ±0.07
VTTS (TV-ordered) 3.62 ±0.20 3.01 ±0.22 3.12 ±0.27
VTTS (VT-ordered) 3.30 ±0.21 3.01 ±0.17 2.35 ±0.22
Table 4: Human evaluation on LRS3. Mean Opinion Scores (MOS) (1-5) with 95% confidence intervals for our VoxCeleb2-trained models evaluated on LRS3. VTTS (TV-ordered) achieves better performance in Intelligibility (3.62) and Synchronization (3.12), while both methods perform equally well in Naturalness (3.01). Ground truth (GT) is an upper bound, while GT (discrete) is an upper bound due to speech quantization. These results demonstrate our model maintains good perceptual quality even on out-of-distribution data.
Method Iterations GT WER ($\downarrow$) GT (discrete) WER ($\downarrow$) WER ($\downarrow$)
TTS 2M 2.6 ±0.1 10.1 ±0.2 17.3$\mathord{+}7.2$
VTTS (TV-ordered) 17.0$\mathord{+}6.9$
VTTS (VT-ordered) $\mathord{+}2.1$
TTS 3M 2.6 ±0.1 10.1 ±0.2 14.7$\mathord{+}4.6$
VTTS (TV-ordered) 14.1$\mathord{+}4.0$
VTTS (VT-ordered) $\mathord{+}2.1$
Table 5: Convergence analysis.   Training iterations comparison shows faster convergence when video modality is used in addition to text. While TTS requires 3M iterations, both VTTS (VT-ordered) and VTTS (TV-ordered) variants achieve comparable or better performance in only 2M iterations. VTTS (VT-ordered) demonstrates superior performance with 12.2% WER and smaller gap (+2.1%) from GT (discrete) WER, showing efficient optimization when leveraging both modalities.
Different aggregation of video representations Table 6 shows results of different strategies for spatial aggregation of video representations, with simple “sum” operation performing the best.
Qualitative results   Figure 5 shows mel-spectrogram comparisons between TTS, GT, and VTTS (VT-ordered). The mel-spectrogram generated by VTTS (VT-ordered) closely resembles GT in terms of temporal structure and speech patterns, particularly in capturing natural pauses and utterance duration. While TTS generates beyond the original duration (445 frames vs GT’s 393 frames) and fails to maintain proper temporal alignment, VTTS (VT-ordered) accurately matches GT’s frame length (393 frames) and successfully captures speech dynamics including pause locations. This demonstrates VTTS’s ability to leverage visual information for generating temporally coherent speech that aligns with the original video timing. The spectral patterns in VTTS (VT-ordered) also show similar energy distributions to GT, particularly in the harmonic structure during speech segments. Analysis of TimeSync for the synchronization is shown in Figures 6 and 7 for the same sample. Influence of modalities    Table 7 shows the impact of ablating individual modalities for VTTS (VT-ordered) model during evaluation. Removing the text modality severely degrades performance, leading to 74.5% WER, while removing the video modality results in 46.4% WER. These results demonstrate that both modalities contribute complementary information, highlighting the importance of our different strategies combining multimodal information.
Attention Average Max Stacking Sum
WER 14.5 13.1 12.4 14.3 12.2
Table 6: Video inputs aggregation. Comparison of different ways to combine video frame representations before inputting into VTTS (TV-ordered) decoder. Sum operation achieves the best WER (12.2%), outperforming more complex strategies like attention (14.5%) and stacking (14.3%). This suggests that simple element-wise operations are sufficient for effective video inputs aggregation.
Method GT WER ($\downarrow$) GT (discrete) WER ($\downarrow$) WER ($\downarrow$)
VTTS (VT-ordered) 2.6 ±0.1 10.1 ±0.2 12.2
    w/o T 74.5
    w/o V 46.4
Table 7: Effect of modality. Word Error Rate (WER) is computed when we systematically remove either text or video modality from the input for our VTTS model during evaluation only. Starting from VTTS (VT-ordered) (12.2% WER), removing text (w/o T) causes significant degradation (74.5% WER), while using only text without video (w/o V) achieves 46.4% WER. These results demonstrate the complementary nature of both modalities, with text providing more content and video significantly contributing to overall speech generation.
Figure 5: Qualitative comparison of log mel-spectrograms. Visualization of generated log mel-spectrograms: Text-to-Speech (TTS, top), Ground Truth (GT, middle), and our Video-Text-to-Speech (VTTS, VT-ordered, bottom). VTTS (VT-ordered) demonstrates better temporal alignment with GT (393 frames) compared to TTS (445 frames), showing the benefit of video conditioning for maintaining correct speech duration. The spectral patterns in VTTS (VT-ordered) also more closely match GT’s energy distribution.
Figure 6: Distribution for TimeSync. We show the difference (left) and absolute difference (right) between ground truth and generated speech phoneme locations (location of the center of the phoneme segment) in time measured in seconds. The ground truth text is used to align it to both ground truth and generated speech. For generated speech we use models TTS, VTTS (VT-ordered) and VTTS (TV-ordered).
Figure 7: Alignment between phonemes in ground truth (GT) speech and generated speech used for TimeSync computation. We show the time in seconds for the phoneme segment centers computed for GT (x-axis) and generated (y-axis) speech for the GT transcription. Dashed gray line is the ideal time synchronization between GT and generated speech. TTS is way out for the proper timing compared to VTTS (VT-ordered) and VTTS (TV-ordered).
4 Related Work
Text-to-Speech Synthesis Text-to-speech (TTS) systems have evolved from early approaches to end-to-end methods [49, 16, 24, 27, 31, 34, 35]. Traditional TTS systems face significant challenges with unseen speaker styles due to substantial enrollment data requirements. While several approaches attempt to address this by extracting speaker representations from speech data [6, 14, 15, 22, 28], obtaining sufficient high-quality utterances remains problematic. Recent studies have incorporated face images for speaker representation [12, 21, 42], aiming to capture visual-acoustic correlations, but often neglect motion-related factors leading to inconsistent voice generation when facial expressions vary. Recent unified architectures for speech-text modeling like VioLA [43] require multi-stage hierarchical processing for EnCodec [11] features, while VOXTLM [25] uses an LM-style approach but relies on HuBERT content tokens losing acoustic and speaker characteristics.
Lip-to-Speech Synthesis Lip-to-speech synthesis aims to reconstruct speech signals from a given face image and silent videos of lips of talking-face, crucial for scenarios with corrupted or missing audio. Early approaches used encoder-decoder architectures with GAN-based training – Lip2Wav [32], End-to-end GAN [30], and VCA-GAN [17] demonstrated success on limited vocabulary datasets, while Lip2Speech [18] extended the GAN framework with multi-task learning for improved content modeling. Recent advances explored discrete token representations through AV-HuBERT [36], with works like ReVISE [13] integrating HiFi-GAN for improved audio generation. In parallel, diffusion models have emerged as a powerful approach for speech generation. Works like DiffWave [20], Grad-TTS [31], and PriorGrad [23] demonstrated effective speech synthesis, leading to LipVoicer [46] which adapted diffusion models for lip-to-speech generation. However, these approaches focus primarily on lip movements, potentially overlooking broader visual dynamics that could improve speech generation. Our work takes a different direction by proposing a novel video-text-to-speech task that leverages complete visual context alongside text input. Rather than using GAN or diffusion-based approaches, we adopt a unified decoder-only transformer architecture inspired by recent successes in LLMs. This enables seamless integration of video, text, and speech modalities for more natural and contextually appropriate speech generation.
5 Conclusion
To the best of our knowledge, we are the first to propose a video-text-to-speech generation framework using a decoder-only transformer architecture. This approach simplifies the multimodal conditional generation of speech while maintaining high-quality output. We demonstrate our approach’s effectiveness by achieving state-of-the-art performance on two challenging datasets: VoxCeleb2 and LRS3 compared to prior approaches that used cropped-lip inputs. These datasets feature diverse speakers, accents, and recording conditions, showcasing our model’s ability to handle real-world scenarios. We formulated a suite of evaluation metrics including Mean Opinion Score for style, synchronization, and content to evaluate naturalness and overall quality of the generated speech. In addition we proposed an automatic metric to assess the quality of alignment between generated and original speech. This multi-faceted evaluation goes beyond traditional metrics to capture nuanced aspects of speech synthesis quality.
6 Acknowledgment
We would like to thank Angelos Katharopoulos for donating the video for the paper, Ruixiang Zhang, Shuangfei Zhai and Russ Webb for fruitful feedback on earlier drafts of the manuscript, Denise Hui for infra and compute support.
Appendix A Ethics Discussion
The advancement of speech technologies brings great potential but also significant ethical challenges that must not be overlooked. While we aim to create techniques that improve conditional speech synthesis for multimodal settings, it is vital to address risks proactively and promote awareness to guide responsible innovation at different levels: from researchers to the end-users. As such, we highlight several key challenges:
  • Dual-use risks   There are always risks of impersonation, voice spoofing attacks, and fake content generation. Safeguarding and watermarking by inserting detectable markers in the generated speech are one of the quickly developing areas to detect the misuse cases.
  • Privacy   We acknowledge the sensitivity of facial and speech data in research and technology development carrying privacy considerations, and thus, we affirm our commitment to protecting individuals’ rights and fostering responsible data usage.
  • Accessibility and inclusivity   While we are working with English-only data for the proof of concept, extending the speech technologies for diverse populations and existing spoken languages should be a top priority in the community.
  • Transparency and accountability   Detailed documentation, limitations, analysis of failure cases, and reproducibility are essential for promoting transparency and informed usage. Responsibility in development and deployment should remain a cornerstone in the community.
Appendix B Limitations
While we made the best effort to tune TTS baseline, there is always a possibility we missed some details. Due to optimization issues when both modalities, video and text, are inputted into the model, we first found best hyper-parameters for our VTTS models so that the models can converge. Later, the same hyper-parameters are used for the TTS baseline by excluding video from the input into the model. However, across all experiments and hyper-parameter tuning we consistently observe that VTTS models outperform TTS models, demonstrating that video brings helpful information for the speech generation.
We did not train larger models ($>$300M parameters), did not use larger datasets ($>$1.5k hours) or pre-trained models, and leave this as a future work.
Appendix C Data, Code, Reproducibility
We made the best effort to use publicly available data and official implementations (e.g. VQ-VAE for video representations). All data we used are under permissive license for research. We do our best to provide all details and steps in the main text and in Appendix. We are in the process of open-sourcing the code and releasing transcriptions PL.v2 for VoxCeleb2 data.
We do not plan to open-source any pre-trained models for sake of privacy, safety and misuse.
Appendix D Video Reconstruction
Although the VQ-VAE model used to extract video representations is pre-trained on the general videos, we found it reconstructs speakers videos with sufficient quality to preserve necessary spatial information. To evaluate video reconstruction quality, we employ the Fréchet Video Distance (FVD) [39], specifically the FVD16 variant that assesses quality over 16-frames window. The FVD scores are computed using an I3D model trained on Kinetics-400, providing a standardized measure of video quality across different temporal scales. The FVD metric is 86.2 at resolution 64x64. Thus, we do not finetune the model further on videos of talking people and use it as is.
Appendix E Evaluation Metrics
Word Error Rate (WER)   We use Whisper-large v2 via open-source code https://github.com/m-bain/whisperX to transcribe generated speech. The latter is compared to the ground truth transcription (PL.v2 is treated as a ground truth for VoxCeleb2) to compute WER.
SyncScore   We are using open-source code for SyncNet from https://github.com/joonson/syncnet_python. During evaluation if generated speech is longer than the video duration then the last video frame is used for the rest of the speech duration.
TimeSync   We use https://github.com/richardbaihe/a3t from [2] to do force alignment between the phoneme sequence of ground truth transcription (PL.v2 is treated as a ground truth for VoxCeleb2) and speech: either generated or original audio. The code is using HMM model from HTK [47] to perform force alignment. This procedure gives us phoneme location in time and phoneme duration for each audio. After, we exclude the silence (“sp”) and its duration from each alignment, see Figure 8. Because every word has several possible phoneme sequences, we use Levenstein distance computation to align sequence of phonemes obtained for the generated and original audio: we consider phonemes to be aligned if they are equal or they can be obtained via substitution operation. Then, we compute the average absolute time difference between the centers of each aligned phoneme segments in generated and original audio, see Figure 9.
TimeSync can be expressed as $\frac{1}{N}\sum_{\phi^{GT}}|t^{model}_{f(\phi^{GT})}-t^{GT}_{\phi^{GT}}|$, where $N$ is the total number of phonemes $\phi^{GT}$ in the ground truth transcriptions obtained for the original audio samples; $t^{GT}_{\phi^{GT}}$ is the segment’s center (in seconds) for the phoneme $\phi^{GT}$ and $t^{GT}_{\phi^{GT}}=(start_{\phi^{GT}}+end_{\phi^{GT}})/2$; $f(\phi^{GT})$ is the phoneme in the phoneme sequence of the generated audio sample which corresponds to ${\phi^{GT}}$ in the alignment; $t^{model}_{f(\phi^{GT})}$ is the segment’s center (in seconds) for the phoneme $f(\phi^{GT})$ and $t^{model}_{f(\phi^{GT})}=(start_{f(\phi^{GT})}+end_{f(\phi^{GT})})/2$; $start$ and $end$ indicate the phoneme’s start and end timestamps in generated or ground truth audio.
Figure 8: TimeSync. Example of the phoneme sequence and its alignment for the ground truth audio before removing silence (“sp”) segments (blue) and after (green).
Figure 9: TimeSync. Example of the phoneme sequence for the ground truth audio (green) and corresponding aligned phoneme sequence for the generated audio (red). TimeSync is computed on these paired (green and red) segments by taking the absolute difference between the segments centers (measure in seconds).
Figure 10: Human evaluation. Task description for the crowd-sourced raters to evaluate intelligibility, naturalness and synchronization of the ground truth or generated speech: speech is overlayed with the video and they are played together for the raters.
Figure 11: Human evaluation. Task description for the crowd-sourced raters to evaluate correspondence between facial expressions and emotions in speech for ground truth and generated speech: speech is overlayed with the video and they are played together for the raters.
Figure 12: Human evaluation. Task description for the crowd-sourced raters to evaluate how close emotions in generated speech follows the ground truth.
Mean Opinion Score (MOS) We use crowd-sourcing to collect subjective ratings to evaluate the intelligibility, naturalness and synchronization of the generated speech. We use the same (randomly sampled) 50 videos from the test set of VoxCeleb2 (or LRS3) [^2] Speakers in the test sets do not overlap with the speakers from the training sets. for each model to generate speech. We then collect around seven ratings per video for each model. Overall, for both VoxCeleb2 and LRS3, we collect 4208 ratings from 387 different raters. The raters were English-speaking and were paid at least the minimum wage.
We present the raters with a generated speech (with volume normalization) overlayed with the original video or original video with original (or reconstructed) speech. We instruct raters to rate how natural speech in the video sounds, how intelligible (e.g. easy to understand) speech is in the video, and how synchronized the speech is with the video on a five-point Likert scale, where 1 corresponds to very unnatural and 5 corresponds to very natural. In Figure 10 we show a screenshot seen by raters. Finally, we compute the MOS with confidence intervals calculated using bootstrap resampling with 10k iterations, providing a reliable estimate of the variability MOS results.
We further instruct raters to evaluate emotional consistency between video and generated speech (’video-speech emotions’) and emotional expressiveness in speech (’speech emotions’) by comparing ground truth and generated audios, see instructions in Figures 11 and 12. MOS results, Table 8, demonstrate the benefit of visual conditioning for emotional expressiveness.
Method video-speech emotions ($\uparrow$) speech emotions ($\uparrow$)
GT 4.62 ±0.07 4.92 ±0.04
GT (discrete) 4.41 ±0.10 4.37 ±0.12
TTS 3.57 ±0.14 3.20 ±0.15
VTTS (TV-streaming) 3.66 ±0.16 3.36 ±0.15
VTTS (TV-ordered) 3.79 ±0.15 3.31 ±0.17
VTTS (VT-ordered) 3.74 ±0.12 3.39 ±0.15
Table 8: Human evaluation of speech-video emotion alignment on VoxCeleb2. Mean Opinion Scores (1-5) with 95% confidence intervals measuring emotional consistency between video and generated speech (’video-speech emotions’) and emotional expressiveness in speech (’speech emotions’). VTTS (TV-ordered) achieves the best performance in video-speech emotion alignment (3.79), while VTTS (VT-ordered) performs best in speech emotion quality (3.39). Both significantly outperform TTS baseline, demonstrating the benefit of visual conditioning for emotional expressiveness. Ground truth (GT) serves as reference, while GT (discrete) shows the upper bound achievable with our approach.
Appendix F Implementation Details
VoxCeleb2 original data has video at 25fps (40ms per frame), or 25Hz, which we use for video representation extraction, while the audio is given at 16kHz and we extract speech representations at 40Hz (25ms per frame).
To select the best hyper-parameters we randomly sampled 2k samples from the training data and use them as the validation data throughout the training. After we find best hyper-parameters on the validation data, we retrain the final models including validation data into training data.
For our VTTS models we stack together speaker embedding, video, text and speech representations. Every modality has prepended begin of sentence representation ($<bos>$) and appended end of sentence representation ($<eos>$). Each modality’s discrete values are mapped to a common dimension $D^{\prime}$ through their respective embedding layers and, optionally, additional linear projections before being fed to the decoder. All our models have $\sim$250M parameters, with $D^{\prime}=768$, 4 heads and 36 transformer layers following the Base architecture from [3]. We follow masking strategy reported in [3]: for every training step with probability $p$ the sample in the minibatch is masked with the mean span of 3 tokens with masking ration of 0.5.
We train final models using the AdamW optimizer with a learning rate of $44e-$ 4, learning rate warmup of $5$k steps, cosine learning rate schedule and gradient clipping of $1.0$. We use dynamic batching to optimize the data packing with total batch size of 16.66 minutes. We train all models till full convergence, with 3M maximum number of steps and with mixed precision training (BF16) on H100 GPUs with 80GB. All models are trained with 8GPUs for 3-5 days.
Appendix G Video-to-Speech
As one of the baselines we trained speech generation model conditioning only on the video input (no text input). The WER for this model is around 100%, and MOS is 1.39 ± 0.10 for intelligibility, 1.60 ± 0.13 for naturalness and 1.49 ± 0.09 for synchronization. The interesting findings about this model are: a) the model is able to generate word $n$grams; b) the model is able to model properly the pauses and reflect the timing when people speaking or being silent.
Appendix H Qualitative Results
In Figures 1416, and 18, we show log mel-spectrogram comparisons between TTS, Ground Truth (GT), and our VTTS (VT-ordered) model across different scenarios. These visualizations include both successful cases where VTTS (VT-ordered) effectively captures temporal dynamics and spectral patterns, and failure cases (Figure  18) that highlight current limitations. Through these examples, we can analyze how video conditioning helps maintain proper speech duration and temporal alignment, while also identifying challenges in generating complex spectral information. Furthermore, to analyze temporal synchronization between generated and ground truth speech, we visualize phoneme-level alignments in Figures 14, 16, and 18. Each plot shows the relationship between phoneme timings in ground truth ($x$-axis) versus generated speech ($y$-axis), where perfect synchronization would follow the diagonal dashed line. The different variants of our model, VTTS (VT-ordered) consistently demonstrate better temporal alignment compared to TTS, as evidenced by their closer distance to the ideal diagonal. This visualization helps quantify how video conditioning helps to maintain proper speech timing and rhythm, with VTTS (VT-ordered) variants showing improved temporal coherence across different examples.
Figure 13: Qualitative comparison of log mel-spectrograms. Visualization of generated log mel-spectrograms: Text-to-Speech (TTS, top), Ground Truth (GT, middle), and our Video-Text-to-Speech (VTTS, bottom). VTTS (VT-ordered) demonstrates better temporal alignment with GT (367 frames) compared to TTS (419 frames), showing the benefit of video conditioning for maintaining correct speech duration. The spectral patterns in VTTS (VT-ordered) also more closely match GT’s energy distribution.
Figure 14: Alignment between phonemes. Temporal alignment visualization for example from Figure 14. The plot compares phoneme timings between ground truth ($x$-axis) and generated speech ($y$-axis). Dashed gray line is the ideal time synchronization between GT and generated speech. TTS is way out for the proper timing compared to VTTS (TV-ordered) and VTTS (TV-ordered).
Figure 15: Qualitative comparison of log mel-spectrograms. Visualization of generated log mel-spectrograms from different methods: Text-to-Speech (TTS, top), Ground Truth (GT, middle), and our Video-Text-to-Speech (VTTS, bottom). VTTS (VT-ordered) demonstrates better temporal alignment with GT (208 frames) compared to TTS (261 frames), showing the benefit of video conditioning for maintaining correct speech duration. The spectral patterns in VTTS (VT-ordered) closely match GT’s harmonic structure and energy distribution, particularly visible in the lower frequency bands (yellow regions). Additionally, VTTS (VT-ordered) accurately captures the temporal dynamics of speech, including pauses and intensity variations, leading to more natural speech generation.
Figure 16: Alignment between phonemes. Temporal alignment visualization for example from Figure 16. The plot compares phoneme timings between ground truth ($x$-axis) and generated speech ($y$-axis). VTTS variants demonstrate superior temporal alignment by following the ideal synchronization line (dashed diagonal) more closely than TTS, which shows significant temporal drift. This example highlights how video conditioning helps maintain proper speech timing.
Figure 17: Failure case analysis of log mel-spectrograms. Visualization of generated log mel-spectrograms from different methods: Text-to-Speech (TTS, top), Ground Truth (GT, middle), and our Video-Text-to-Speech (VTTS, bottom). While VTTS (VT-ordered) maintains better temporal alignment with GT (165 frames vs TTS’s 176 frames), both VTTS (VT-ordered) and TTS struggle to accurately capture GT’s harmonic structure and energy distribution. Despite having video conditioning, VTTS (VT-ordered) shows degraded spectral quality particularly in the mid-frequency ranges, though it still preserves some temporal speech dynamics like pauses. This example highlights current limitations in generating complex spectral patterns.
Figure 18: Alignment between phonemes. Temporal alignment visualization for failure case corresponding to Figure 18. Plot shows phoneme timing comparison between ground truth ($x$-axis) and generated speech ($y$-axis). While VTTS variants maintain better alignment than TTS, all models show deviation from ideal synchronization (dashed diagonal), particularly in later segments, illustrating challenges in maintaining temporal coherence for complex speech patterns.